OpenStudy (anonymous):

An online company called “Animals Dressed as Humans” is making matching owner/dog socks for the upcoming winter. A total of 19 humans and dogs purchased socks. They sold a total of 60 individual socks. How many humans and how many dogs purchased socks to keep their feet warm? Part 1: Set up the problem. Part 2: Explain why you set it up that way. Part 3: Provide the solution.

hero (hero):

h = humans d = dogs Set up system of equations h + d = 19 2h + 4d = 60 I chose to set up systems of equations because it is easier to keep up with each variable and also, when using x and y as variables, I can graph both equations to find the correct (x,y) coordinates.

OpenStudy (anonymous):

Thank you (:

hero (hero):

You will still need to do part III which is to find the solution. Basically, you have to find h and d.

OpenStudy (anonymous):

Allow x = humans and y = dogs. Then, x + y = 19 represents the total number of orders that were placed. 2x + 4y = 60 represents the total number of socks that were sold (2 because humans have two feet and 4 because dogs have four feet). Then, solve the first equation for x, to get x = 19 - y, and substitue this into the second equation for x. You get 2(19 - y) + 4y = 38 - 2y + 4y = 38 + 2y = 60. Subtract 38 from both sides to get 2y = 22 and divide both sides by 2 to get y = 11. Finally substitute this value back into the first equation for y, to get x + 11 = 19, and by subtracting 11 from both sides, x = 8. Therefore, 8 humans and 11 dogs purchased socks to keep their feet warm. To check, multiply 8(2) to get 16 and 11(4) to get 44, and 16 + 44 = 60! :)

OpenStudy (anonymous):

did that help :)

OpenStudy (anonymous):

Oops. Lol I already submitted it. But thanks anyways. & yeah it would've